OpenStudy (chris215):

The following is a geometric sequence. -5, 15, 45, -225

OpenStudy (chris215):

true or false

OpenStudy (mathmath333):

\(\large\tt \color{black}{ common~~ ratio=\dfrac{15}{-5}=-3 \neq\dfrac{45}{15}=3\neq\dfrac{-225}{45}=-5}\)

OpenStudy (mathmath333):

false